(7/12) + (11/12) = 3/2
 (7/12) + (11/12) = 3/2
 (7/12) + (11/12) = 3/2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement isTrue
Consider Left Hand Side (LHS),
Sum of like fractions = (7/12) + (11/12)

= (7 + 11)/12

= 18/12
Divide both numerator and denominator by 6, we get,

= 3/2 Right Hand Side (RHS) = 3/2 By comparing LHS and RHS,

LHS = RHS
3/2 = 3/2



Thanks